Write the standard form of the equation of the circle with the given conditions.

1a. Center (9, -10) and r=5

1b. Center (2, 0) and r=square root of 2
1
Expert's answer
2012-12-06T09:53:10-0500
The common form of the equation of the circle is

(x - x0)² + (y - y0)² = r²,

where (x0, y0) is the center point of the circle and r is its radius. Therefore,
&
1a. Center (9, -10)& and r=5

(x - 9)² + (y - (-10))² = 5²,

or

(x - 9)² + (y + 10)² = 25.
&
1b. Center (2, 0) and r=square root of 2

(x - 2)² + (y - 0)² = (√2)²,

or

(x - 2)² + y² = 2.